OpenStudy (anonymous):

The average of the test scores of a class of n students is 82, and the average of the test scores of another class of p students is 89. If the average of the combined classes is 84, what is the value of p/n?

OpenStudy (kropot72):

The total marks of class #1 = 82n The total marks of class #2 = 89p The total marks of classes #1 and #2 combined = 84(n + p) Now the following equation can be formed: 84(n + p) = 82n + 89p 84n + 84p = 82n + 89p ............(1) The required value of p/n can now be found from equation (1)

OpenStudy (kropot72):

@sakigirl Do you want me to help you further?

OpenStudy (anonymous):

@kropot72 is it 4/25?

OpenStudy (kropot72):

How did you arrive at that result?

OpenStudy (anonymous):

Well, I got 2n=5p and I divided both sides by 2, and then 5, and put it in a fraction equaling p and n @kropot72

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Your equation 2n = 5p is correct. Now divide both sides by 5 and get \[p=\frac{2}{5}n\] Now divide both sides by n to get the solution.

OpenStudy (anonymous):

@kropot72 Thank you! I have another question. Would you be willing to answer that too?

OpenStudy (kropot72):

You're welcome :). Please post your other question and I'll try to help.